# Implicit High-Order Moment Tensor Estimation and Learning Latent Variable Models

### v1

@@ 1.2 Main Result: General Implicit Moment Tensor Estimation
Proposition 1.7 gives an efficient method to approximate higher moment tensors of the form $\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$ for non-negative $w_{i}$’s, where the order $t$ is large, so long as there are “reasonable” unbiased estimators for these moments that can be efficiently sampled. This result broadly generalizes the work by Li and Liu [LL22], who introduced the idea of implicit moment tensor estimation in the context of clustering mixtures of separated spherical Gaussians and, more broadly, separated mixtures of translates of any known Poincare distribution $D$. Roughly speaking, using the terminology introduced above, they gave an efficient algorithm for implicit moment tensor estimation only for the specific STC $\mathcal{S}_{t}$ arising in their clustering application—namely, corresponding to the particular expression they used for the adjusted polynomials that come up in their clustering model.
The key conceptual contribution of our work is the realization that the computational task of implicit moment tensor estimation can be solved efficiently in a vastly more general setting, and thus applied to several other learning problems. This answers the main open question posed in [LL22]; see, e.g., [Liu24]. To achieve this goal, we introduce our general framework involving the notion of an STC (Definition 1.6)—which is new to this work. It turns out that this type of sequence of tensors arises in a variety of probabilistic models as Fourier or moment tensors. Learning these moment tensors is thus a powerful tool and, as we have shown (Section 1.1), leads to more efficient learning algorithms for these models. Beyond its generality, this broader view leads to quantitative improvements even for the specific clustering application of [LL22], allowing us to cluster mixtures of spherical Gaussians under the optimal separation of $\sqrt{\log(k)}$; see Theorem 1.3 and Remark 4.2.
@@ 1.3 Technical Overview
The naive method to approximate $\langle T(\mathcal{X}),M_{m}\rangle$ is as follows: Since $M_{m}$ is the expectation of the output $\mathcal{S}_{m}(\mathcal{D}_{m})$, where $\mathcal{D}_{m}$ was assumed to be efficiently samplable, we can approximate $M_{m}$ by drawing sufficiently many samples from $\mathcal{D}_{m}$, computing the random output tensors (obtained from $\mathcal{S}_{m}$ on each sample), and averaging them. Similarly, we can average the output tensors of $\mathcal{F}_{m}$, obtained via drawing sufficiently many samples from the relevant input distribution $\mathcal{D}^{\prime}$, to get an approximation to $T(\mathcal{X})$. Computing the inner product of these approximations yields an approximation to the desired answer.
@@
Statistical Challenges.
We start with the statistical aspect. Note that both $T(\mathcal{X})$ and $M_{m}$ are order-$m$ tensors in $d$ dimensions, i.e., they lie in a $d^{m}$-dimensional space. Thus, in the absence of further structural properties, in order to approximate the means of the sequential tensor computations $\mathcal{S}_{m}$ and $\mathcal{F}_{m}$ to $\ell_{2}$-error $\epsilon$, one would need roughly $\Omega(d^{m}(V/\epsilon^{2}))$ samples, where $V$ is the covariance upper bound. Unfortunately, this sample upper bound is prohibitively large for our learning applications—indeed, in the underlying applications, we need to take $m=\omega(1)$. [^3] Concretely, for learning mixtures of $k$ spherical Gaussians one would need $m=\Omega(\log(k))$. For learning sums of ReLU activations, we would need to select $m=\mathrm{poly}(1/\epsilon)$, where $\epsilon$ is the desired $L_{2}$-error. Finally, for learning $k$-mixtures of linear regressions with Gaussian noise of standard deviation $\sigma$, one would need $m=\Omega(\log(1/\epsilon)/\sigma^{2})$.
In order to overcome this obstacle, we need a method of reducing the dimension of the underlying space. We can achieve this by considering moment tensors $M_{t}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$ of different orders. Suppose, for example, that we knew exactly the moment tensor $M_{2m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes 2m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}\otimes v_{i}^{\otimes m}$. When viewed as a $d^{m}\times d^{m}$ matrix, $M_{2m}$ will be rank-$k$ with image (span of its column vectors) $W$ spanned by the $v_{i}^{\otimes m}$’s. Thus, if we already knew $M_{2m}$, we could compute its image $W$ and then note that $M_{m}$ lies in $W$. Then, instead of computing the full output tensors of the sequential tensor computations $\mathcal{S}_{m}$ and $\mathcal{F}_{m}$—which are objects lying in $(\mathbb{R}^{d})^{\otimes m}$—we could compute their projections onto $W$. Since $\dim(W)\leq k$ (as $M_{2m}$ is rank-$k$ by definition), by standard concentration arguments it follows that $O(kV/\epsilon^{2})$ samples suffice to compute these projections to $\ell_{2}$-error $\epsilon$.
Computational Challenges.
@@
In this subsection, we sketch how our main result is leveraged to derive our learning applications. We start with the problem of learning sums of ReLUs, followed by our density estimation algorithms for MLRs and GMMs, and concluding with our application on clustering GMMs.
PAC Learning Sums of ReLUs
Recall that the objective is to approximate a function of the form $F(x)=\sum_{i=1}^{k}w_{i}\mathrm{ReLU}(v_{i}\cdot x)$, where $w_{i}\geq 0$, given access to random labeled examples $(x,F(x))$ with $x\sim N(0,I)$. We proceed by attempting to learn the low degree part of its Fourier transform. In particular, we have that
$F(x)=\sum_{n=0}^{\infty}\langle T_{n},H_{n}(x)\rangle$
where $H_{n}$ is the normalized Hermite tensor (Definition 2.2) and $T_{n}=\mathbf{E}[F(G)H_{n}(G)]$. Our approach relies on approximating $F$ by the sum of the first $\mathrm{poly}(1/\epsilon)$ many terms in the sum above. To do this, we note that the tensors $T_{n}$ are moment tensors of approximately the kind desired. Namely, it can be shown (see Lemma 4.8) that
$T_{n}=c_{n}\sum_{i=1}^{k}w_{i}v_{i}^{\otimes n}\;,$
for some known constants $c_{n}$ (which are critically bounded away from 00 when $n$ is even). Given this, we can use Proposition 1.7 directly to approximate $\langle T_{n},H_{n}(x)\rangle$ in time roughly $\mathrm{poly}(k,d)2^{O(n)}$. Doing this for all $n$ up to degree $\mathrm{poly}(1/\epsilon)$ yields the desired approximation (proof of Theorem 4.7).
@@ 1.4 Prior and Related Work
A line of prior work on learning MLRs focused on the parameter estimation problem. Specifically, a sequence of paper (see, e.g., [ZJD16, LL18, KC19] and references therein) analyzed non-convex methods, including expectation maximization and alternating minimization. These works established local convergence guarantees: Given a sufficiently accurate solution (warm start), these non-convex methods can efficiently boost this to a solution with arbitrarily high accuracy. The focus of our algorithmic results in this section is to provide such a warm start, which captures the complexity of the problem. We note that the local convergence result of [LL18] applies for the noiseless case, while the result of [KC19] can handle non-trivial regression noise when the weights of the unknown mixture are known.
The prior works most closely related to ours are [LL18, CLS20, DK20]. The work of [LL18] studied the noiseless setting (corresponding to $\sigma=0$) and provided an algorithm with sample complexity and running time scaling exponentially with $k$. Subsequently, [CLS20] gave an algorithm whose sample complexity and running time scales with $\exp(\tilde{O}(k^{1/2}))$, i.e., sub-exponential in $k$. In more detail, for the noisy case, when $\sigma=O(\epsilon)$ and the weights are uniform, the algorithm of [CLS20] has sample complexity and runtime of the form $\mathrm{poly}(dk/(\epsilon\Delta))\,(k/\epsilon)^{\tilde{O}(k^{1/2}/\Delta^{2})}$, where $\Delta$ is the minimum pairwise separation between the regressors. The fastest previously known algorithm [DK20] for both parameter and density estimation has sample and computational complexity scaling quasi-polynomially in $k$. Specifically, for density estimation the latter algorithm has complexity $\mathrm{poly}(d,(k/\epsilon)^{\log^{2}(k)})$.
@@
In the related task of parameter estimation, the goal is to approximate the parameters of the components (i.e., mixture weights and component means). This task requires further assumptions, so that it is solvable with polynomial sample complexity (even information-theoretically). The typical assumption in the literature is that pairwise separation between the component means is bounded below by some parameter $\Delta>0$. A long line of work since the late 90s steadily improved the separation requirement [Das99, AK01, VW02, AM05, KSV08, BV08, RV17, HL18, KSS18, DKS18, LL22]. The state-of-the-art prior to this paper is the work of [LL22] that gave a $\mathrm{poly}(d,k,1/\epsilon)$ time algorithm under near-optimal separation of $\Theta((\log(k/\epsilon))^{1/2+c})$. Finally, we note that the assumption on spherical component covariances is crucial for the latter result (and our Theorem 1.3). Namely, a departure from sphericity leads to information-computation tradeoffs: [DKPZ23] gave evidence that quasi-polynomial runtime is inherent for any $\mathrm{polylog}(k)$ separation, even if the Gaussian components have the same bounded covariance.
Learning One-hidden-layer ReLU Networks
@@
There has been a recent explosion of research on provable algorithms for learning neural networks in various settings, see, e.g., [JSA15, SJA16, DFS16, ZLJ16, ZSJ+17, GLM18, GKLW19, BJW19, GKKT17, GK19, VW19, DKKZ20, GGJ+20, DK20, CKM21, CGKM22, CDG+23, DK24] for some works on the topic. Many of these works focused on parameter learning—the problem of recovering the weight matrix of the data generating neural network. PAC learning of simple classes of networks has been studied as well  [GKKT17, GK19, VW19, DKKZ20, CGKM22, CDG+23, DK24].
@@
The work of [GLM18] studies the parameter learning of positive linear combinations of ReLUs under the Gaussian distribution. It is shown in [GLM18] that the parameters can be efficiently approximated, if the weight matrix is full-rank with bounded condition number. The complexity of their algorithm scales polynomially with the condition number. [BJW19, GKLW19] obtained efficient parameter learners for vector-valued depth-$2$ ReLU networks under the Gaussian distribution. Similarly, the algorithms in these works have sample complexity and running time scaling polynomially with the condition number.
@@
In contrast to parameter estimation, PAC learning these function classes does not require any assumptions on the structure of the weight matrix. The PAC learning problem for this class is information-theoretically solvable with polynomially many samples. The question is whether a computationally efficient algorithm exists. Until recently, the problem of PAC learning positive linear combinations of ReLUs For the task of positive linear combinations of ReLUs studied in this work, [DKKZ20] gave a learner with sample complexity $\mathrm{poly}(dk/\epsilon)$ and computational complexity $\mathrm{poly}(dk/\epsilon)+(k/\epsilon)^{O(k^{2})}$. This runtime bound was improved to $\mathrm{poly}(d)(k/\epsilon)^{O(\log^{2}(k))}$ in [DK20], which was the state-of-the-art prior to our work.
@@ 1.5 Organization
The structure of the paper is as follows: In Section 2, we provide the necessary definitions and technical facts. In Section 3, we prove our main result. Section 4 presents our learning algorithms for sums of ReLUs, mixtures of spherical Gaussians, and mixtures of linear regressions.
@@ 2 Preliminaries
Let $V$ be an inner product space. If $A$ and $B$ are elements of $V^{\otimes t}$ for some $t\in\mathbb{Z}_{+}$, then we use $\langle A,B\rangle$ to denote the inner product of $A$ and $B$ induced by the inner product on $V$. We also use $\|A\|_{2}=\langle A,A\rangle^{1/2}$ for the corresponding $\ell_{2}$-norm.
A tensor $A$ of order $t$ and dimensions $(d_{i})_{i\in[t]}$ is a multilinear map defined by a $t$-dimensional array with real entries $A_{\alpha}$, where $\alpha=(\alpha_{1},\ldots,\alpha_{t})$ with $\alpha_{i}\in[d_{i}]$. The covariance of an order-$t$, dimension $d$ tensor-valued random variable $T\in(\mathbb{R}^{d})^{\otimes t}$ is defined as the tensor in $(\mathbb{R}^{d})^{\otimes 2t}$ defined as $\mathbf{Cov}[T]:=\mathbf{E}[T\otimes T]-\mathbf{E}[T]\otimes\mathbf{E}[T]$, where the expectation operation is applied entry-wise. Similarly, the corresponding “second moment” is the tensor $\mathbf{E}[T\otimes T]$. For an order-$t$, dimension-$d$ tensor $W$, consider the real-valued random variable $\langle W,T\rangle$ and its corresponding variance, $\mathbf{Var}[\langle W,T\rangle]$. We say that $\mathbf{Cov}[T]$ is bounded above by $C$, for some $C>0$, if $\max_{\|W\|_{2}=1}\mathbf{Var}[\langle W,T\rangle]\leq C$, i.e., if the variance of $T$ in any (normalized) direction is at most $C$. The analogous definition applies for the second moment $\mathbf{E}[T\otimes T]$.
@@
Before we describe and analyze our algorithm, we introduce some necessary technical tools.

### v2

@@ Discussion
Proposition 1.8 gives an efficient method to approximate higher moment tensors of the form $\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$ for non-negative $w_{i}$’s, where the order $t$ is large, so long as there are “reasonable” unbiased estimators for these moments that can be efficiently sampled. This result broadly generalizes the work by Li and Liu [LL22], who introduced the idea of implicit moment tensor estimation in the context of clustering mixtures of separated spherical Gaussians and, more broadly, separated mixtures of translates of any known Poincare distribution $D$. Roughly speaking, using the terminology introduced above they gave an efficient algorithm for implicit moment tensor estimation only for the specific tensors arising in their clustering application—namely, the parameter moment tensors computed in a specific way using their so-called adjusted polynomials.
The key conceptual contribution of our work is the realization that the computational task of implicit moment tensor estimation can be solved efficiently in a vastly more general setting, and thus applied to several other learning problems. This answers the main open question posed in [LL22]; see, e.g., [Liu24].
It turns out that the kind of sequence of tensors $M_{t}=\sum_{i}w_{i}v_{i}^{\otimes t}$, that our technique requires, arises in a variety of probabilistic models as Fourier or moment tensors. We also generalize the way in which these tensors can be computed from the expectation of an adjusted-polynomial to the output of any Sequential Tensor Computation (Definition 3.1)—which is new to this work. Learning these moment tensors is thus a powerful tool and, as we have shown (Section 1.1), leads to more efficient learning algorithms for these models. Beyond its generality, this broader view leads to quantitative improvements even for the specific clustering application of [LL22], allowing us to cluster mixtures of spherical Gaussians under the optimal separation of $\sqrt{\log(k)}$; see Theorem 1.3 and Remark 4.2.
@@ 1.3 Technical Overview
The naive method to approximate $\langle T,M_{m}\rangle$ is as follows: Since $M_{m}$ is the expectation of $\mathcal{S}_{m}$, we can approximate $M_{m}$ by drawing sufficiently many samples from this distribution and averaging them. We can approximate $T$ similarly by sampling from $\mathcal{F}_{m}$, and computing the inner product of these approximations yields an approximation to the desired answer.
@@
Statistical Challenges
We start with the statistical aspect. Note that both $T$ and $M_{m}$ are order-$m$ tensors in $d$ dimensions, i.e., they lie in a $d^{m}$-dimensional space. Thus, in the absence of further structural properties, in order to approximate them to error $\epsilon$ by sampling from $\mathcal{S}_{m}$ and $\mathcal{F}_{m}$, one would need roughly $\Omega(d^{m}(V/\epsilon^{2}))$ samples, where $V$ is the covariance upper bound. Unfortunately, this sample upper bound is prohibitively large for our learning applications—indeed, in the underlying applications, we need to take $m=\omega(1)$. [^2] Concretely, for learning mixtures of $k$ spherical Gaussians one would need $m=\Omega(\log(k))$. For learning sums of ReLU activations, we would need to select $m=\mathrm{poly}(1/\epsilon)$, where $\epsilon$ is the desired $L_{2}$-error. Finally, for learning $k$-mixtures of linear regressions with Gaussian noise of standard deviation $\sigma$, one would need $m=\Omega(\log(1/\epsilon)/\sigma^{2})$.
In order to overcome this obstacle, we need a method of reducing the dimension of the underlying space. We can achieve this by considering moment tensors $M_{t}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$ of different orders. Suppose, for example, that we knew exactly the moment tensor $M_{2m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes 2m}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}\otimes v_{i}^{\otimes m}$. When viewed as a $d^{m}\times d^{m}$ matrix, $M_{2m}$ will be rank-$k$ with image (span of its column vectors) $W$ spanned by the $v_{i}^{\otimes m}$’s. Thus, if we already knew $M_{2m}$, we could compute its image $W$ and then note that $M_{m}$ lies in $W$. Then, instead of computing the full output tensors of the $\mathcal{S}_{m}$ and $\mathcal{F}_{m}$—which are objects lying in $(\mathbb{R}^{d})^{\otimes m}$—we could compute their projections onto $W$. Since $\dim(W)\leq k$ (as $M_{2m}$ is rank-$k$ by definition), by standard concentration arguments it follows that $O(kV/\epsilon^{2})$ samples suffice to compute these projections to $\ell_{2}$-error $\epsilon$.
Computational Challenges
@@
In this subsection, we sketch how Proposition 1.8 is leveraged to obtain our learning applications. We start with the problem of learning sums of non-linear activations, followed by our density estimation algorithms for MLRs and GMMs, and concluding with our application on clustering GMMs.
PAC Learning Sums of Non-linear Activations
Recall that the goal of this task is to approximate a function of the form $F(x)=\sum_{i=1}^{k}w_{i}\sigma(v_{i}\cdot x)$, where $w_{i}\geq 0$ and $\sigma$ is an appropriate (non-linear) activation function, given access to random labeled examples $(x,F(x))$ with $x\sim N(0,I)$. We proceed by attempting to learn the low degree part of the Fourier transform of $F$. In particular, we have that
$F(x)=\sum_{m=0}^{\infty}\langle T_{m},H_{m}(x)\rangle$
where $H_{m}$ is the normalized Hermite tensor (Definition 2.2) and $T_{m}=\mathbf{E}[F(G)H_{m}(G)]$. Our approach relies on approximating $F$ by the sum of the first $n$ many terms in the sum above. To do this, we note that the tensors $T_{m}$ are moment tensors of approximately the kind desired. Namely, it can be shown (see Lemma 4.12) that
$T_{m}=c_{\sigma,m}\sum_{i=1}^{k}w_{i}v_{i}^{\otimes m}\;,$
for some known constants $c_{\sigma,m}$ (which are critically bounded away from 00 when $m$ is even). Given this, we can use Proposition 1.8 directly to approximate $\langle T_{m},H_{m}(x)\rangle$ in time roughly $\mathrm{poly}(kd/\epsilon)2^{O(m)}$. Doing this for all $m$ up to degree $n$ yields the desired approximation (see proof of Theorem 4.7 for the detailed argument).
@@ 1.4 Prior and Related Work
Since their introduction [DeV89, JJ94], mixtures of linear regressions (MLRs) have been broadly studied as a generative model for supervised data, and have found applications to various problems, including trajectory clustering and phase retrieval; see [Che21] for additional discussion and references. A line of algorithmic work focused on the parameter estimation problem. Specifically, a sequence of papers (see, e.g., [ZJD16, LL18, KC19] and references therein) analyzed non-convex methods, including expectation maximization and alternating minimization. These works established local convergence guarantees: Given a sufficiently accurate solution (warm start), these non-convex methods can efficiently boost this to a solution with arbitrarily high accuracy. The focus of our algorithmic results is to provide such a warm start, which captures the complexity of the problem. We note that the local convergence result of [LL18] applies for the noiseless case, while the result of [KC19] can handle non-trivial regression noise when the weights of the unknown mixture are known.
The prior works most closely related to ours are [LL18, CLS20, DK20]. The work of [LL18] studied the noiseless setting (corresponding to $\sigma=0$) and provided a parameter estimation algorithm with sample complexity and running time scaling exponentially with $k$. Subsequently, the work of [CLS20] gave a sub-exponential time parameter estimation algorithm for both the noiseless case and the case where $\sigma$ is small, namely $\sigma=O(\epsilon)$. Specifically, if the weights are uniform and $\sigma=O(\epsilon)$, their algorithm has sample and computational complexity $\mathrm{poly}(dk/(\epsilon\Delta))\,(k/\epsilon)^{\tilde{O}(k^{1/2}/\Delta^{2})}$, where $\Delta$ is the minimum pairwise separation between the regressors. The fastest previously known algorithm [DK20] for both parameter and density estimation has sample and computational complexity scaling quasi-polynomially in $k$. Specifically, for density estimation (the task we focus on in this paper) the algorithm of  [DK20] has complexity $\mathrm{poly}(d,(k/\epsilon)^{\log^{2}(k)})$.
@@
In the related task of parameter estimation, the goal is to approximate the parameters of the components (i.e., mixture weights and component means). This task requires further assumptions, so that it is solvable with polynomial sample complexity (even information-theoretically). The typical assumption in the literature is that pairwise separation between the component means is bounded below by some parameter $\Delta>0$. A long line of work since the late 90s steadily improved the separation requirement [Das99, AK01, VW02, AM05, KSV08, BV08, RV17, HL18, KSS18, DKS18, LL22].
A related line of work studies the parameter estimation problem in a smoothed setting, e.g., under the assumption that the means are linearly independent and the corresponding matrix has bounded condition number. Specifically, [HK13] gave a polynomial-time learner for such instances. Interestingly, under such assumptions, access to constant-order moments suffices. We emphasize that in the general case that we focus on (i.e., without such assumptions) access to super-constant degree moments is necessary.
The state-of-the-art algorithm for parameter estimation (in the general case) prior to this paper is the work of [LL22]: they gave a $\mathrm{poly}(d,k,1/\epsilon)$ time algorithm under near-optimal separation of $\Theta((\log(k/\epsilon))^{1/2+c})$. We note that the assumption on spherical component covariances is crucial for the latter result (and our Theorem 1.3). Namely, a departure from sphericity leads to information-computation tradeoffs: [DKPZ23] gave evidence that quasi-polynomial runtime is inherent for any $\mathrm{polylog}(k)$ separation, even if the Gaussian components have the same bounded covariance.
Learning One-hidden-layer Networks
@@
Over the past decade, we have witnessed an explosion of research activity on provable algorithms for learning neural networks in various settings, see, e.g., [JSA15, SJA16, DFS16, ZLJ16, ZSJ+17, GLM18, GKLW19, BJW19, GKKT17, GK19, VW19, DKKZ20, GGJ+20, DK20, CKM21, CGKM22, CDG+23, DK24, CN24a] for some works on the topic. Many of these works focused on parameter learning—the problem of recovering the weight matrix of the data generating neural network. PAC learning of simple classes of networks has been studied as well  [GKKT17, GK19, VW19, DKKZ20, CGKM22, CDG+23, DK24].
@@
The work of [GLM18] studies the parameter learning of positive linear combinations of ReLUs and other non-linear activations under the Gaussian distribution. It is shown in [GLM18] that the parameters can be efficiently approximated, if the weight matrix is full-rank with bounded condition number. The complexity of their algorithm scales polynomially with the condition number. [BJW19, GKLW19] obtained efficient parameter learners for vector-valued depth-$2$ ReLU networks under the Gaussian distribution. Similarly, the algorithms in these works have sample complexity and running time scaling polynomially with the condition number.
@@
In contrast to parameter estimation, PAC learning does not require any assumptions on the structure of the weight matrix. The PAC learning problem for one-hidden-layer networks is information-theoretically solvable with polynomially many samples. The question is whether a computationally efficient algorithm exists. For the task of positive linear combinations of ReLUs studied in this work, [DKKZ20] gave a learner with computational complexity $\mathrm{poly}(dk/\epsilon)+(k/\epsilon)^{O(k^{2})}$. This runtime bound was improved to $\mathrm{poly}(d)(k/\epsilon)^{O(\log^{2}(k))}$ in [DK20], which was the state-of-the-art prior to our work.
@@ 1.5 Organization
The structure of the paper is as follows: In Section 2, we provide the necessary definitions and technical facts. In Section 3, we prove our main algorithmic result on implicit tensor estimation. Section 4 presents our learning algorithms for sums of ReLUs, mixtures of spherical Gaussians, and mixtures of linear regressions.
@@ 2 Preliminaries
Fix some $t\in\mathbb{Z}_{+}$. For vector spaces $V_{i}$, $i\in[t]$, an order-$t$ tensor on $\bigotimes_{i=1}^{t}V_{i}$ is simply an element $A\in\bigotimes_{i=1}^{t}V_{i}$. If $A,B\in\bigotimes_{i=1}^{t}V_{i}$, for some inner product spaces $V_{i}$, $i\in[t]$, then we use $\langle A,B\rangle$ to denote the inner product of $A$ and $B$ induced by the inner products on $V_{i}$. Namely, using the unique inner product structure on $\bigotimes_{i=1}^{t}V_{i}$ so that $\langle v_{1}\otimes\cdots\otimes v_{t},w_{1}\otimes\cdots\otimes w_{t}\rangle=\prod_{i=1}^{t}\langle v_{i},w_{i}\rangle$ for all vectors $v_{i},w_{i}\in V_{i}$. For the special case that $V_{i}=\mathbb{R}^{d_{i}}$, $d_{i}\in\mathbb{Z}_{+}$, a tensor $A$ is defined by a $t$-dimensional array with real entries $A_{\alpha}$, where $\alpha=(\alpha_{1},\ldots,\alpha_{t})$ with $\alpha_{i}\in[d_{i}]$. For tensors $A,B$ of order $t$ with the same dimensions, the inner product $\langle A,B\rangle$ is the entrywise inner product of the corresponding arrays. We also use $\|A\|_{2}=\langle A,A\rangle^{1/2}$ for the corresponding $\ell_{2}$-norm.
The covariance of an order-$t$, dimension $d$ tensor-valued random variable $T\in(\mathbb{R}^{d})^{\otimes t}$ is defined as the tensor in $(\mathbb{R}^{d})^{\otimes 2t}$ defined as $\mathbf{Cov}[T]:=\mathbf{E}[T\otimes T]-\mathbf{E}[T]\otimes\mathbf{E}[T]$, where the expectation operation is applied entry-wise. Similarly, the corresponding “second moment” is the tensor $\mathbf{E}[T\otimes T]$. For an order-$t$, dimension-$d$ tensor $W$, consider the real-valued random variable $\langle W,T\rangle$ and its corresponding variance, $\mathbf{Var}[\langle W,T\rangle]$. We say that $\mathbf{Cov}[T]$ is bounded above by $C$, for some $C>0$, if $\max_{\|W\|_{2}=1}\mathbf{Var}[\langle W,T\rangle]\leq C$, i.e., if the variance of $T$ in any (normalized) direction is at most $C$. The analogous definition applies for the second moment $\mathbf{E}[T\otimes T]$.
@@
The structure of this section is as follows: Section 3.1 presents our formal framework and the detailed statement of Proposition 1.8. In Section 3.2, we establish some technical results that are necessary for our implicit tensor estimation algorithm and its analysis. Finally, in Section 3.3 we give the pseudocode of our algorithm and prove its correctness.
3.1 Formal Framework and Statement of Main Result
To state our result, we first need to formalize what we mean by a “nice” arithmetic circuit to compute our tensors. The key point is that this method of computation must be compatible with our method for implicitly representing the subspaces $W_{t}$ (described in Section 1.3.1): namely, repeatedly tensoring with $\mathbb{R}^{d}$ and then applying some linear transformation back to $\mathbb{R}^{k}$. Compatibly with this, we can add two elements of a single $W_{t}$ or multiply one by a scalar. We can also take the tensor product of a tensor in $W_{t}$ with an element of $\mathbb{R}^{d}$ and map the result to $W_{t+1}$. Unfortunately, we cannot easily perform other operations like take the tensor product of two elements of $W_{t}$ and find its image in $W_{2t}$. The notion of a Sequential Tensor Computation below amounts to an arithmetic circuit that performs only operations compatible with these computations.
Definition 3.1 (Sequential Tensor Computation).
  • (i)
    Multiplication of a tensor, vector, or scalar by a scalar.
  • (ii)
  • (iii)
We note that an STC is a general way to produce higher-order tensors that is compatible with the kinds of computations that we need to perform on them (see Lemmas 3.7 and 3.8) in the context of learning applications.
Our main algorithmic result (i.e., the formal version of Proposition 1.8) is the following:
Proposition 3.2 (Main Algorithmic Result on Implicit Tensor Computation).
Let $d,k\in\mathbb{Z}_{+}$, $w_{i}\in\mathbb{R}_{\geq 0}$ and $v_{i}\in\mathbb{R}^{d}$ for all $i\in[k]$. Consider the sequence of dimension-$d$ tensors $\{M_{t}\}$ with $M_{t}$ order-$t$, for $t\in\mathbb{Z}_{+}$, defined by $M_{t}=\sum_{i=1}^{k}w_{i}v_{i}^{\otimes t}$.
Let $m\in\mathbb{Z}_{+}$. Suppose that for all positive integers $t\leq 2m$, with $t$ even or equal to $m$, there exists an STC $\mathcal{S}_{t}$ of order-$t$ and size at most $S$ with the following property: when given as input $\mathcal{I}$ a sample drawn from some efficiently samplable distribution $\mathcal{D}_{t}$, the output tensor $\mathcal{S}_{t}(\mathcal{I})$, $\mathcal{I}\sim\mathcal{D}_{t}$, has mean $M_{t}$ and covariance bounded above by $V$, for some $V>0$. Let $\mathcal{F}_{m}$ be another STC of order-$m$ and size at most $S$, whose input is partitioned as $(\mathcal{X},\mathcal{Y})$, with the following properties: when $\mathcal{X}\sim X$ and $\mathcal{Y}\sim\mathcal{D}^{\prime}$, where $\mathcal{D^{\prime}}$ is efficiently samplable, the tensor $\mathcal{F}_{m}(\mathcal{X},\mathcal{Y})$ has expectation over $\mathcal{Y}\sim\mathcal{D^{\prime}}$
$T(\mathcal{X}):=\mathbf{E}_{\mathcal{Y}\sim\mathcal{D^{\prime}}}[\mathcal{F}_{m}(\mathcal{X},\mathcal{Y})]$ (1)
and second moment over both $\mathcal{X}$ and $\mathcal{Y}$ bounded by $V$.
$\mathbf{E}_{\mathcal{X}\sim X}[(\mathcal{A}-\langle T(\mathcal{X}),M_{m}\rangle)^{2}]\leq\frac{\mathrm{poly}\left(k,m,d,V,1/\tau,1+\mathop{\textstyle\sum}_{i=1}^{k}w_{i}\right)\left(1+\max_{i\in[k]}\|v_{i}\|_{2}\right)^{2m}}{N^{1/2}}\;.$ (2)
Discussion
In words, we show that, with high probability over the samples drawn from $\mathcal{D}_{t}$ and $\mathcal{D}^{\prime}$, the expected squared error of our estimator over a random choice of $\mathcal{X}\sim X$ is bounded as shown in (2) above. In particular, our algorithm can be thought of as taking samples from $\mathcal{D}_{t}$ and $\mathcal{D}^{\prime}$ as input and returning an evaluator for a function of $\mathcal{X}$ that with high probability is close in $L_{2}$ to the function $\langle T(\mathcal{X}),M_{m}\rangle$.
Remark 3.3.
It turns out that this kind of generalization described above is necessary for our applications—rather than simply finding the inner product of $M_{m}$ with a particular tensor $T$. This is because, for example, we might want to compute $\langle H_{m}(x),M_{m}\rangle$ for the Hermite tensor $H_{m}$ evaluated at a random $x$. While we could simply let $T=H_{m}(x)$ for some given $x$, that would run into the problem that $H_{m}(x)$ is itself quite large—so a naive application of our result would give large error. However, as the second moment of $H_{m}(x)$ is bounded, our result says that the mean squared error over a random choice of $x$ is small.
A couple of additional remarks are in order. First, we did not attempt to optimize the polynomial dependence on the relevant parameters, as this would not affect the qualitative aspects of our learning theory applications. Second, the dependence on the failure probability $\tau$ can be made logarithmic, i.e., $\log(1/\tau)$, by running the algorithm $O(\log(1/\tau))$ times and taking a median; such a dependence is not required in our applications.